OpenStudy (anonymous):

sum of 2 angles is 180. find measures of both if one of these angles is 28 degrees less than 3 times the measure of the other.

OpenStudy (anonymous):

Let angle 1= x Let angle 2 = y Therefore x+y = 180 From the second part: 3x-28=y or: 3x-y=28 Can you do the rest?

OpenStudy (anonymous):

wouldnt it be 3x-28=y or -3x +y= -28 instead ??????

OpenStudy (anonymous):

I said that. All I did was rearrange it.